OpenStudy (anonymous):

a number added to 8 times the numbers reciprocal is 6. find the numbers

OpenStudy (asnaseer):

Do you know how to express this as an equation?

OpenStudy (asnaseer):

cal the unknown number 'x'. then what the question states is:\[x+\frac{8}{x}=6\]do you know how to solve this?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

first multiply both sides of the equation by 'x' to get rid of the fraction. this gives us:\[x^2+8=6x\]which can be re-arranged to:\[x^2-6x+8=0\]this is a standard quadratic equation. hopefully you know how to solve this?

OpenStudy (anonymous):

no

OpenStudy (asnaseer):

do you know how to factorise this expression?

OpenStudy (anonymous):

mmm no

OpenStudy (anonymous):

can u please tell me the answer

OpenStudy (anonymous):

i got it is 2 and 4

OpenStudy (asnaseer):

ok - looks like you are trying to solve a problem without first learning the foundation behind it. It might be better for you to first review any notes that your teacher may have given you and then try again to solve this. The aim of OpenStudy is to help students understand how to solve problems - not just give them the answer.

OpenStudy (anonymous):

4 and 2

OpenStudy (asnaseer):

@ale5 - how did you get 2 and 4 if you do not know how to solve this?

OpenStudy (anonymous):

i factor 8 and it give me -2 and -4 and if you add htem is -6 then i solve for x

OpenStudy (asnaseer):

ok - that is the correct approach. you confused me above by saying that ou didn't know how to factorise.

OpenStudy (asnaseer):

I guess what you were trying to say is that you know how to factorise but just couldn't "see" what the factors were for this problem?